OpenStudy (anonymous):

solve: 2-3(x-4)=8+3(2-x)

OpenStudy (unklerhaukus):

\[2-3(x-4)=8+3(2-x)\] expand the brackets \[2-3\times(x)-3\times(-4)=8+3\times(2)+3\times(-x)\]

OpenStudy (unklerhaukus):

now simplify